Calculate the amount of heat released (kJ) in the complete oxidation of 6.95 g of Al metal at standard condition. 

4 Al(s) + 3 O2(g) → 2 Al2O3(s)      ΔH° = -3352 kJ


1
Expert's answer
2021-02-04T07:07:46-0500

Atomic mass of Al is 27 g/mole

6.95g of aluminium =(6.95/27)=0.25moles of Al

According to given equation 4 moles Aluminium produce 3352kj heat during oxidation,

0.25 moles Al produce (3352*0.25/4)=209.5kj